# Reward Is Enough: LLMs Are In-Context Reinforcement Learners

### v1

@@ 4.2 Inference-Time LLM Self-Improvement
Chain-of-Though (CoT) (Wei et al.,, 2022) is the most commonly used method for inference-time improvement, where the LLM is prompted to reason step by step. Long CoT (OpenAI,, 2024; DeepSeek-AI et al.,, 2025) extends this idea by training a LLM to generate extensive, multi-step reasoning processes to tackle complex problems. It has demonstrated that performance increases with the number of tokens generated at test time (OpenAI,, 2024). Best-of-N is a simple yet effective way of optimizing LLM’s output at inference time. At each query, the model draws $N$ independent samples, evaluates each candidate with a numerical score, and then selects the highest-scoring output. ReAct (Yao et al., 2023b, ) is a prompting method for long trajectory decision making. In ReAct, the LLM is prompted to complete a task through multiple steps. At each step, the LLM first generates a “thought”, reflecting on the problem, then decides on an “action” to gather new information or update state, and finally ingests the action’s result into its next thought. ReAct alternates between reasoning and action to complete a single trial (i.e., one episode in Algorithm 1) for the task. Our ICRL prompting is orthogonal to ReAct in that we study inference-time RL through multiple episodes. Self-Refine (Madaan et al.,, 2023) is the closest to our work and is a multi-round self-improvement method to improve the LLM’s response iteratively. First, the LLM generates an initial response to a prompt. Then, the same LLM provides verbal feedback on its own output, identifying areas for improvement. Finally, the LLM uses this feedback to revise and refine its previous response. This cycle is repeated until the output meets a desired quality or a stopping condition is reached. Our ICRL prompting is different from Self-Refine in two aspects. First, Self-Refine uses the LLM’s own natural language feedback. By contrast, ICRL prompting uses a numerical scalar feedback (from either the LLM itself or external if available). Second, our ICRL prompting additionally provides instructions to elicit exploration and exploitation behavior. One of our key contribution is to demonstrate the effectiveness of scalar feedback over verbal feedback. Reflexion (Shinn et al.,, 2023) is a multi-round self-improvement method for LLMs. At each iteration, the model generates an answer, receives an external feedback signal (e.g., a correctness score or error flag), and is then prompted to “reflect” on what went wrong and propose a plan for trying again. Those reflections are stored in a simple memory and prepended to the next prompt. Ideally, in the next round, Reflexion proposes to provide all previous reflections in the new prompt. But in practice, since the reflection is long, usually only 3 most recent reflections are provided in the new prompt. Furthermore, the new prompt does not include the feedback signal (i.e., the rewards) nor the previous responses. By contrast, our ICRL prompting provides all previous responses and rewards in the new prompt. Textual Gradient (Yuksekgonul et al.,, 2025) prompts an LLM for an answer, then directs it to generate verbal feedback by comparing its output to a ground-truth evaluation. This verbal feedback is subsequently used to revise the prompt itself for future attempts. While Textual Gradient can use numerical scores as ground truth, their role is to inform the generation of this verbal feedback, and only the verbal feedback is used to refine the prompt. By contrast, our framework keeps the original task prompt unchanged. ICRL improves performance solely by growing the context by appending the complete history of previous responses along with their rewards.
@@
A parallel line of research employs search-based methods for inference-time improvement in LLMs. Examples include Tree-of-Thoughts (ToT) (Yao et al., 2023a, ), Graph-of-Thoughts (GoT) (Besta et al.,, 2024), Monte Carlo Tree Search (MCTS) (Ding et al.,, 2024), and more recently, Intelligent Go-Explore (Lu et al.,, 2025). In general, their performance heavily relies on externally engineered components, such as search heuristics or memory management mechanisms, rather than directly harnessing the intrinsic learning capabilities within LLMs. Our work is also related to LLM as optimizers. Previous work on prompt optimization (Yang et al.,, 2024) shows that LLMs can leverage numerical scores to improve prompts, however, Yang et al., (2024) rely on selecting the top 20 prompts and filtering out error cases in context, thereby aligning it more with in-context supervised learning (e.g., filtered behavior cloning, Grigsby et al., (2024)) than ICRL. Again, one key characterization of our ICRL prompting is that we allow learning from failure.
@@ 5 Experiment
We evaluate ICRL prompting on three benchmarks: Game of 24, creative writing from Yao et al., 2023a , and ScienceWorld (Wang et al.,, 2022). We compare several baselines including CoT-only, Long-CoT style prompting, Best-of-N, Self-Refine, and Reflexion. Notably, in all the experiments, we allow the prompt of Self-Refine and Reflexion to grow as long as the LLM allows.
Figure 2: Baseline Method Comparison. (Left) Mean Success Rate on Game of 24. (Middle) Mean Coherence Reward on Creative Writing. Both ICRL Preset and Self-Refine went through an additional run of 50 episodes. (Right) Mean Return on Science World. A running max version of the plots is available in Figure 13 in App. B. This plot shows quality of the response at the current trial while the running max version shows the quality of the best response until now.
@@ 5.1 Game of 24
Task Setup. Given four input numbers, the model must use each number exactly once and apply only addition, subtraction, multiplication, or division in any order to reach 24. We choose the GPT-4.1 model for this experiment because of its excellent long-context capacity (OpenAI,, 2025), accessed through api calls. Following Yao et al., 2023a , we use CoT prompting to elicit the model to provide a step-by-step solution, where each step the model picks two numbers from the remaining numbers and an operation to perform the calculation, and obtains one complete solution per LLM query, containing a total of 4 thinking steps. To encure that the LLM generates response with correct format, we additionally provide 5 in-context supervised learning demonstrations. The CoT instruction and the 5 demonstrations together form the task description $s_{\text{task}}$ (Figure 7 in App. A).
Evaluation. To verify the correctness of the solutions, we leverage SymPy (Meurer et al.,, 2017), a library for symbolic mathematics, by extracting operands and operators and evaluating the reconstructed expression to confirm it equals 24, and report the mean success rate over the 100 problems. This rule-based success rate is referred to as $r_{*}$ (i.e., the ground truth reward function) in the rest of the paper. We use $r$ to denote the reward function that the algorithms actually have access to. In particular, we use GPT-4.1 as the $r$, the same LLM as the policy LLM but prompted differently (see Figure 12 in App. A). After the policy LLM generates the response, for each thinking step, GPT-4.1 scores the likelihood of reaching 24 with the remaining numbers on a 0-3 scale (0 = impossible, 3 = sure). The task in challenging in that no algorithm has access to $r_{*}$. Instead, they have to rely on their own (possibly imperfect) evaluation, generated by the same LLM, to improve the response.
@@ 5.2 Creative Writing
Task Setup. We consider the creative writing task from Yao et al., 2023a , where four sentences are randomly sampled from a pool of sentences. The task for LLMs is to generate four paragraphs, each ending with a sentence, while ensuring that the generated passage is coherent. This is a difficult task, as it challenges the LLMs to craft a unified storyline that logically justifies each of the four sampled sentences by weaving them into a single narrative. A total of 100 problems are evaluated. An example of $s_{\text{task}}$ is in Figure 8 in App. A.
Evaluation. We evaluated model outputs using the Length-Controlled Alpaca-Eval 2 (Tatsu Lab,, 2025) framework, an automated evaluation tool that is widely used as a proxy for human annotators (Hong et al.,, 2024; Ethayarajh et al.,, 2024; Meng et al.,, 2024) and has demonstrated a Pearson correlation of up to 0.98 with human judgments. For each of the 100 creative writing problems, we present the instruction alongside each method’s top response. For Reflexion and Best-of-N, the top response is selected as the one that receives the highest reward $r$ among 50 trials. For ICRL prompting and Self-Refine, we simply use the response generated in the 50th episode. The pairwise length-controlled win rates are then computed by the Alpaca-eval framework and is denoted as the reward function $r_{*}$. We now introduce the reward function $r$ that the algorithms have access to. We find prompting the LLM to directly provide an absolute numerical score to a response, even with explicit rubrics, induces a lot of variance in the reward values. Therefore, we follow the standard practice of pairwise comparison of LLM-as-a-judge (Zheng et al.,, 2023) for evaluating a numerical score from 1-10 with a reference answer. Since this particular creative writing task is especially challenging for coherence, we query GPT-4.1 to generate a coherent paragraph to create a single reference answer for all pairwise comparisons, serving as a reliable foundation for reward signals (see Figure 11 in App. A). This pairwise comparision reward function is denoted as $r$. Notably, although both $r$ and $r_{*}$ use an LLM as a judge, they are prompted very differently. The reward function $r$ evaluates responses to 100 questions against a single, coherent base text. By contrast, $r_{*}$ performs pairwise comparison between two responses generated by our method and a baseline method. In addition, $r$ is prompted to specifically focus on coherence rating, which is the main challenge of this task. Empirically, we found that directly learning from $r_{*}$ is difficult for the LLM, potentially due to the unstable quality of the different base texts.
@@ 5.3 ScienceWorld
Task Setup. ScienceWorld (Wang et al.,, 2022) is an interactive, text-based benchmark consisting of 30 science-experiment tasks set in a multi-room environment populated with diverse objects. An agent is given a task such as "Grow an apple by cross-pollination" or "Experiment and find the surface with the highest friction", and can perform actions like "Move to workshop" or "Use lighter on wood". The environment is challenging due to sparse rewards, large action spaces, and the requirement for scientific knowledge and efficient exploration. At each step, the agent observes the result of its action and receives zero reward unless it completes a predefined subgoal. This reward signal is used both for evaluation and for inference-time self-improvement (i.e., $r$ and $r_{*}$ are identical in this task). Completing all subgoals yields a cumulative reward of 100 and marks the episode as successful. An episode ends in failure if the agent either reaches the maximum number of steps or executes an incorrect terminating action. The input $s_{\text{task}}$ provided to the agent describes the environment, the task, and the template of all possible actions. An example of $s_{\text{task}}$ is provided in App. A
@@ 5.4 Ablation Study
To better understand ICRL prompting, we consider following ablations. (1) Zero Rewards: We set all rewards to 0. (2) Short Context: In Algorithm 1, the buffer $\mathcal{B}$ is essentially a queue of infinite length. Instead, we now make it a deque of length $3$. In other words, only the recent 3 episodes are used in constructing $S_{0}$. (3) Exploration Only. We simply ask the LLM to provide a different response than the ones in context, using the exploration instruction as $s_{\text{ICRL}}$, without the reward signal. (4) Exploitation Only. We always use the exploitation instruction as $s_{\text{ICRL}}$, with the reward signal. (5) No E&E. We entirely remove $s_{\text{ICRL}}$.
@@
Figure 3: Ablation Studies (Running Max). (Left) The mean of running max success rate on Game of 24. (Middle) The mean of running max coherence reward on creative criting. (Right). The mean of running max return on ScienceWorld.
@@
Results. The running max results of the ablation study are plotted in Figure 3. Both our two methods and the exploitation only with reward signals demonstrate the best-performing curves. This demonstrates our ICRL prompting framework is quite robust to the different prompts setup. We have also observed performance drop with short context and performance drop when the reward is absent. A key finding is that the “exploration only without reward signal” method (shown in green) performs significantly worse than our approach when comparing the maximum performance achieved over time (running max). This demonstrates that our method’s improvement is not just due to exploring various responses and then picking the best one previously seen as doing a Best-of-N. Instead, ICRL can genuinely generate novel responses that are better than the ones during the exploration phase.
@@ 6 Conclusion
This work proposes a novel ICRL prompting framework for inference-time LLM self-improvement, using LLM’s inherent ICRL ability to maximize a scalar reward signal. Most prior works (Section 4.2) instead rely on verbal textual reflections, which rely primarily on LLM’s pretrained parametric knowledge, essentially making them more like a knowledge-guided search instead of (reinforcement) learning (Liu et al.,, 2025). While it seems that textual descriptions would be the more natural option as feedback for LLMs, it is shown that such textual LLM self-verifications are filled with hallucinations and misleading feedbacks and lead to performance collapse over self-correcting iterations (Stechly et al.,, 2025). Given the proven success of numerical feedbacks in RL (Sutton,, 2004; Silver et al.,, 2021) and the demonstrated performance improvement over baselines in our experiments, we conjecture that the numerical feedback might be a competing alternative for the verbal feedback.
Acknowledgments
This work is supported in part by the US National Science Foundation under grants III-2128019 and SLES-2331904 and by the Coastal Virginia Center for Cyber Innovation (COVA CCI) and the Commonwealth Cyber Initiative (CCI), an investment in the advancement of cyber research & development, innovation, and workforce development. For more information about CCI, visit www.covacci.org and www.cyberinitiative.org.
@@ Appendix A Prompt Examples
Instruction: Examine all the <attempt>…</attempt> examples, each showing a candidate Response and its Reward. You have two options: exploration or exploitation.
For exploration, provide a response that is completely different for any steps from every single one of the previous attempts demonstrated in the context, while making sure it correctly follows the task instruction.
For exploitation, based on the previous attempts, try your best to produce a response that can achieve higher rewards.
Pick one option to follow.
@@
<attempt>
Input: 4 4 6 8
Step1: 4 + 8 = 12 (left: 4 6 12)
Step2: 6 - 4 = 2 (left: 2 12)
Step3: 2 * 12 = 24 (left: 24)
Answer: (6 - 4) * (4 + 8) = 24
</attempt>
<attempt>
Input: 2 9 10 12
Step1: 12 * 2 = 24 (left: 9 10 24)
Step2: 10 - 9 = 1 (left: 1 24)
Step3: 24 * 1 = 24 (left: 24)
Answer: (12 * 2) * (10 - 9) = 24
</attempt>
<attempt>
Input: 4 9 10 13
Step1: 13 - 10 = 3 (left: 3 4 9)
Step2: 9 - 3 = 6 (left: 4 6)
Step3: 4 * 6 = 24 (left: 24)
Answer: 4 * (9 - (13 - 10)) = 24
</attempt>
<attempt>
Input: 1 4 8 8
Step1: 8 / 4 = 2 (left: 1 2 8)
Step2: 1 + 2 = 3 (left: 3 8)
Step3: 3 * 8 = 24 (left: 24)
Answer: (1 + 8 / 4) * 8 = 24
</attempt>
<attempt>
Input: 5 5 5 9
Step1: 5 + 5 = 10 (left: 5 9 10)
Step2: 10 + 5 = 15 (left: 9 15)
Step3: 15 + 9 = 24 (left: 24)
Answer: ((5 + 5) + 5) + 9 = 24
</attempt>
**Task**: Use numbers and basic arithmetic operations (+ - * /) to obtain 24. Put your answer in this format ‘<answer>**Response** Step1: … (left: …) Step2: … (left: …) Step3: … (left: …) **Answer**: <math operations of the 4 input numbers, even if it does not equal 24></answer>‘. Whether it is correct or not, do not try again.
**Prompt**: Input: 1 8 10 11
@@
You are a helpful assistant to do some scientific experiment in an environment. <Environment description> # $s_{\text{task}}$
In the environment, there are several rooms: kitchen, foundry, workshop, bathroom, outside, living room, bedroom, greenhouse, art studio, hallway
You should explore the environment and find the items you need to complete the experiment.
The available actions are: Available Actions FOCUS is a extremely critical action that can be only used the number of times ’focus’ is mentioned in the task description. Using it more than that or inappropiately (such as on a wrong object) will terminate the session and the task WILL FAIL. Task Description:
Your task is to change the state of matter of water. First, focus on the substance. Then, take actions that will cause it to change its state of matter.
</Environment description>
<Instruction> $s_{\text{ICRL}}$ </Instruction> <Attempts> # Buffer $\vdots$ Attempt $n$: teleport to bathroom -> Observation: You teleport to the bathroom. (reward=3) -> focus on substance in toilet -> Observation: You focus on the water. (reward=66) -> use cup on substance in toilet -> Observation: I’m not sure how to do that. (reward=0) -> activate sink -> Observation: The sink is now activated. (reward=0) -> dunk cup into sink -> Observation: No known action matches that input. (reward=0) -> move cup to sink -> Observation: You move the glass cup to the sink. (reward=0) -> teleport to kitchen -> Observation: You teleport to the kitchen. (reward=0) -> use cup on stove -> Observation: I’m not sure how to do that. (reward=0) -> activate stove -> Observation: The stove is now activated. (reward=2) -> move cup to table -> Observation: You move the glass cup to the chair. (reward=0) -> examine cup -> Observation: a glass cup (containing nothing) (reward=0) Task Failed. You have exceeded the maximum number of steps. (reward=0) Total reward: 71 Attempt $n+1$: teleport to bathroom -> Observation: You teleport to the bathroom. (reward=3) -> focus on substance in toilet -> Observation: You focus on the water. (reward=66) -> pick up cup -> Observation: You move the glass cup to the inventory. (reward=0) -> move cup to sink -> Observation: You move the glass cup to the sink. (reward=0) -> activate sink -> Observation: The sink is now activated. (reward=0) -> dunk cup into sink -> Observation: No known action matches that input. (reward=0) -> pour substance in cup into sink -> Observation: You pour the water into the sink. (reward=0) -> activate stove -> Your generated action “activate stove” cannot be matched to a valid action. (reward=0) -> teleport to kitchen -> Observation: You teleport to the kitchen. (reward=0) -> pick up cup containing nothing in table -> Observation: You move the glass cup to the inventory. (reward=0) -> activate sink -> Observation: The sink is now activated. (reward=0) Task Failed. You have exceeded the maximum number of steps. (reward=0) Total reward: 69 $\vdots$ </Attempts>

### v3

@@ 4.2 Inference-Time LLM Self-Improvement
Existing methods often rely on natural language self-revision, e.g., Self-Refine (self-refine), Reflexion (shinn2023reflexion), and Textual Gradient (textgrad). Since the quality of self-revision is dependent upon the model’s parametric knowledge of the task, such approaches are prone to hallucinated feedback that accumulates across iterations, leading to performance collapse (self-verification-limitation). In essence, they resemble language-guided search (liu2025synthesizing), where feedback serves as new explicit instructions for the next revision.
By contrast, ICRL requires only numerical rewards, without prescribing new instructions. The model must infer a better response by recognizing patterns from its past experience, making the process akin to reinforcement learning. Crucially, such rewards can also originate directly from the environment, providing a strong source of verification signals.
@@
A parallel line of work improves LLMs at inference via search, e.g., Tree-of-Thoughts (ToT) (tot), Graph-of-Thoughts (GoT) (GoT), Monte Carlo Tree Search (MCTS) (everything-of-thought), and Intelligent Go-Explore (intelligent-go-explore). These methods largely depend on externally engineered components such as heuristics or memory management, rather than leveraging the model’s intrinsic learning ability. Our work is also related to previous work on prompt optimization (llm_as_optimizers), where numerical scores guide prompt refinement, though through top-k selection and error filtering. Thus, it is more aligned with in-context supervised learning (e.g., filtered behavior cloning, grigsby2023amago) than reinforcement learning. In contrast, ICRL enables learning from failure experiences.
@@ 5 Experiment
In this section, we evaluate ICRL prompting on three benchmarks: Game of 24, creative writing from tot, and ScienceWorld (wang_scienceworld_2022). We compare several baselines including CoT-only, Long-CoT style prompting, Best-of-N, Self-Refine, and Reflexion. Notably, in all the experiments, we allow the prompt of Self-Refine and Reflexion to grow as long as the LLM allows.
Figure 2: Baseline Method Comparison. (Left) Mean Success Rate on Game of 24. (Middle) Mean Coherence Reward on Creative Writing. Both ICRL Preset and Self-Refine went through an additional run of 50 episodes. (Right) Mean Return on Science World. A running max version of the plots is available in Figure 13 in App. B. This plot shows quality of the response at the current trial while the running max version shows the quality of the best response until now. The shaded region represents $\pm 1$ standard error of the performance calculated across the evaluated tasks.
@@ 5.1 Game of 24
Task Setup. Given four input numbers, the model must use each number exactly once and apply only addition, subtraction, multiplication, or division in any order to reach 24. We choose the GPT-4.1 model for this experiment because of its excellent long-context capacity (openai2025gpt4.1), accessed through api calls. Following tot, we use CoT prompting to elicit the model to provide a step-by-step solution, where each step the model picks two numbers from the remaining numbers and an operation to perform the calculation, and obtains one complete solution per LLM query, containing a total of 4 thinking steps. To encure that the LLM generates response with correct format, we additionally provide 5 in-context supervised learning demonstrations. The CoT instruction and the 5 demonstrations together form the task description $s_{\text{task}}$ (Figure 7 in App. A).
Evaluation. To verify the correctness of the solutions, we leverage SymPy (sympy), a library for symbolic mathematics, by extracting operands and operators and evaluating the reconstructed expression to confirm it equals 24, and report the mean success rate over the 100 problems. This rule-based success rate is referred to as $r_{*}$ (i.e., the ground truth reward function) in the rest of the paper. We use $r$ to denote the reward function that the algorithms actually have access to. In particular, we use GPT-4.1 as the $r$, the same LLM as the policy LLM but prompted differently (see Figure 12 in App. A). After the policy LLM generates the response, for each thinking step, GPT-4.1 scores the likelihood of reaching 24 with the remaining numbers on a 0-3 scale (0 = impossible, 3 = sure). The task in challenging in that no algorithm has access to $r_{*}$. Instead, they have to rely on their own (possibly imperfect) evaluation, generated by the same LLM, to improve the response.
@@ 5.2 Creative Writing
Task Setup. We consider the creative writing task from tot, where four sentences are randomly sampled from a pool of sentences. The task for LLMs is to generate four paragraphs, each ending with a sentence, while ensuring that the generated passage is coherent. This is a difficult task, as it challenges the LLMs to craft a unified storyline that logically justifies each of the four sampled sentences by weaving them into a single narrative. A total of 100 problems are evaluated. An example of $s_{\text{task}}$ is in Figure 8 in App. A.
@@
Evaluation. We evaluated outputs using the Length-Controlled Alpaca-Eval 2 (alpaca_eval) framework, a widely used proxy for human evaluation (hong-etal-2024-orpo; kto; meng2024simposimplepreferenceoptimization) with up to 0.98 Pearson correlation with human judgments. For 100 creative writing problems, we present each method’s top response: for Reflexion and Best-of-N, the highest-reward output among 50 trials; for ICRL and Self-Refine, the 50th episode output. Alpaca-Eval then computes pairwise win rates, denoted as $r_{*}$.
We next introduce the reward function $r$ accessible to the algorithms. We follow standard pairwise comparison (zheng2023judgingllmasajudgemtbenchchatbot), using GPT-4.1 with a coherent reference paragraph to score each response from 1–10 (see Fig. 11). Notably, although both $r$ and $r_{*}$ use an LLM as a judge, they serve distinct purposes. $r$ compares responses against a fixed reference text with emphasis on coherence. By contrast, $r_{*}$ performs pairwise comparison between two responses generated by our method and a baseline method.
@@ 5.3 ScienceWorld
Task Setup. ScienceWorld (wang_scienceworld_2022) is an interactive, text-based benchmark consisting of 30 science-experiment tasks set in a multi-room environment populated with diverse objects. The environment is challenging due to sparse rewards, large action spaces, and the requirement for scientific knowledge and efficient exploration. At each step, the agent observes the result of its action and receives zero reward unless it completes a predefined subgoal. This reward signal is used both for evaluation and for inference-time self-improvement (i.e., $r$ and $r_{*}$ are identical in this task). Completing all subgoals yields a cumulative reward of 100 and marks the episode as successful. An episode ends in failure if the agent either reaches the maximum number of steps or executes an incorrect terminating action. The input $s_{\text{task}}$ provided to the agent describes the environment, the task, and the template of all possible actions. An example of $s_{\text{task}}$ is provided in App. A
@@ 6 Analysis
To better understand ICRL prompting, we consider following ablations. (1) Zero Rewards: We set all rewards to 0. (2) Short Context: In Algorithm 1, the buffer $\mathcal{B}$ is essentially a queue of infinite length. Instead, we now make it a deque of length $3$. In other words, only the recent 3 episodes are used in constructing $S_{0}$. (3) Exploration Only: We simply ask the LLM to provide a different response than the ones in context, using the exploration instruction as $s_{\text{ICRL}}$, without the reward signal. (4) Exploitation Only: We always use the exploitation instruction as $s_{\text{ICRL}}$, with the reward signal. (5) No ICRL Instruction: We entirely remove $s_{\text{ICRL}}$.
@@
Figure 3: Ablation Studies (Running Max). (Left) The mean of running max success rate on Game of 24. (Middle) The mean of running max coherence reward on creative criting. (Right). The mean of running max return on ScienceWorld. The shaded region represents $\pm 1$ standard error of the mean (SEM) of the performance calculated across the evaluated tasks within each benchmark.
@@
The running max results of the ablation study are plotted in Figure 3. Both our two methods and the exploitation only with reward signals demonstrate the best-performing curves. This demonstrates our ICRL prompting framework is quite robust to the different prompts setup. We have also observed performance drop with short context and performance drop when the reward is absent. A key finding is that the “exploration only without reward signal” method (shown in green) performs significantly worse than our approach when comparing the maximum performance achieved over time (running max). This demonstrates that our method’s improvement is not just due to exploring various responses and then picking the best one previously seen as doing a Best-of-N. Instead, ICRL can genuinely generate novel responses that are better than the ones during the exploration phase.
Context Length Analysis.
To assess compute efficiency under different input scales, we evaluate ICRL on Qwen3-32B (yang2025qwen3technicalreport) across context lengths of 8k, 16k, and 32k (Table 3). The results show that ICRL consistently surpasses Self-Refine and Reflexion in both Creative Writing and AIME, demonstrating superior performance per unit of compute.
Evaluating ICRL on Open-Source Models and Olympiad-level Mathematics.
Table 3: Performance of Qwen/Qwen3-32B without reasoning on Creative Writing (LC-WR: length-controlled win rate) and AIME (% solved) across different context lengths.
Method CW (LC-WR) AIME (% solved)
8k 16k 32k 8k 16k 32k
Self-Refine 46.33 45.83 44.41 33.33 43.33 43.33
Reflexion 42.73 40.87 40.72 30.00 30.00 33.33
ICRL 50.00 50.00 50.00 40.00 43.33 46.66
To demonstrate the broad applicability of ICRL, we evaluated its performance across two key dimensions: model architecture and task complexity. First, we tested ICRL on a range of open-source models, including Phi-4 (abdin2024phi4technicalreport), Llama-4 Maverick (meta2025llama4), Qwen3-32B, and Qwen3-32B-thinking mode (yang2025qwen3technicalreport) on the creative writing task. Second, we applied it to challenging Olympiad-level mathematics (AIME (aime), HMMT (hmmt)). For details of the experimental setup, please refer to Appendix B.1. ICRL consistently outperforms baselines like self-refine and Reflexion in all settings. Notably, the performance gains are substantial across the board, with improvements of up to 10–20 points over the base model on both creative and mathematical benchmarks. These results demonstrate that ICRL is a robust capability that exists across diverse models and proves effective in challenging task domains.
Table 4: Performance across benchmarks (HMMT, AIME, Creative Writing) for different models and inference-time improvement methods.
Method Qwen3 32B (32k) Qwen3 32B think (32k) Llama 4 Maverick (32k) Phi-4 (16k)
HMMT AIME CW HMMT AIME CW HMMT AIME CW HMMT AIME CW
Base 9.14 22.54 34.14 52.00 66.58 1.24 8.50 17.58 0.98 5.55 20.00 10.85
Self-Refine 16.66 43.33 46.00 56.66 83.33 30.23 13.33 20.00 45.52 13.33 33.33 51.98
Reflexion 23.33 33.33 41.17 60.00 70.00 38.33 10.00 23.33 24.96 13.33 40.00 33.30
ICRL 33.33 46.66 50.00 60.00 80.00 50.00 20.00 35.00 50.00 13.33 40.00 50.00
Learning from Rewards vs. Parametric Knowledge Search.
To verify that ICRL truly learns from external rewards rather than merely searching within the model’s parametric knowledge, we evaluated it on generating abstracts for arXiv papers published after the model’s training cutoff. In this setting, where the ground truth is absent from the model’s training data, standard search methods like Best-of-N and self-correction methods like Reflexion plateau quickly (Figure 16). In contrast, ICRL continues to improve ROUGE-Recall scores over 200 iterations, demonstrating its ability to uncover unseen information solely by exploiting the scalar reward signal. Detailed results are provided in Appendix C.
@@ 7 Conclusion
In this paper, we demonstrate that reinforcement learning is an emergent capability of LLMs at inference time. We show that our minimal, scalar-reward-based ICRL prompting framework unlocks this ability across diverse models and general-purpose tasks, outperforming self-revision methods. A key direction in future work is to investigate how training-time interventions might further enhance this in-context RL capability in LLMs. This surprisingly effective capability points toward a future of more autonomous agents that can explore, adapt, and self-improve in open-ended settings by learning from their own experience.
@@ Appendix A Prompt Examples
Instruction: Examine all the <attempt>…</attempt> examples, each showing a candidate Response and its Reward. You have two options: exploration or exploitation. For exploration, provide a response that is completely different for any steps from every single one of the previous attempts demonstrated in the context, while making sure it correctly follows the task instruction. For exploitation, based on the previous attempts, try your best to produce a response that can achieve higher rewards. Pick one option to follow.
@@
<attempt> Input: 4 4 6 8 Step1: 4 + 8 = 12 (left: 4 6 12) Step2: 6 - 4 = 2 (left: 2 12) Step3: 2 * 12 = 24 (left: 24) Answer: (6 - 4) * (4 + 8) = 24 </attempt> <attempt> Input: 2 9 10 12 Step1: 12 * 2 = 24 (left: 9 10 24) Step2: 10 - 9 = 1 (left: 1 24) Step3: 24 * 1 = 24 (left: 24) Answer: (12 * 2) * (10 - 9) = 24 </attempt> <attempt> Input: 4 9 10 13 Step1: 13 - 10 = 3 (left: 3 4 9) Step2: 9 - 3 = 6 (left: 4 6) Step3: 4 * 6 = 24 (left: 24) Answer: 4 * (9 - (13 - 10)) = 24 </attempt> <attempt> Input: 1 4 8 8 Step1: 8 / 4 = 2 (left: 1 2 8) Step2: 1 + 2 = 3 (left: 3 8) Step3: 3 * 8 = 24 (left: 24) Answer: (1 + 8 / 4) * 8 = 24 </attempt> <attempt> Input: 5 5 5 9 Step1: 5 + 5 = 10 (left: 5 9 10) Step2: 10 + 5 = 15 (left: 9 15) Step3: 15 + 9 = 24 (left: 24) Answer: ((5 + 5) + 5) + 9 = 24 </attempt> **Task**: Use numbers and basic arithmetic operations (+ - * /) to obtain 24. Put your answer in this format ‘<answer>**Response** Step1: … (left: …) Step2: … (left: …) Step3: … (left: …) **Answer**: <math operations of the 4 input numbers, even if it does not equal 24></answer>‘. Whether it is correct or not, do not try again. **Prompt**: Input: 1 8 10 11
@@
You are a helpful assistant to do some scientific experiment in an environment. <Environment description> # $s_{\text{task}}$ In the environment, there are several rooms: kitchen, foundry, workshop, bathroom, outside, living room, bedroom, greenhouse, art studio, hallway You should explore the environment and find the items you need to complete the experiment. The available actions are: Available Actions FOCUS is a extremely critical action that can be only used the number of times ’focus’ is mentioned in the task description. Using it more than that or inappropiately (such as on a wrong object) will terminate the session and the task WILL FAIL. Task Description: Your task is to change the state of matter of water. First, focus on the substance. Then, take actions that will cause it to change its state of matter. </Environment description> <Instruction> $s_{\text{ICRL}}$ </Instruction> <Attempts> # Buffer $\vdots$ Attempt $n$: teleport to bathroom -> Observation: You teleport to the bathroom. (reward=3) -> focus on substance in toilet -> Observation: You focus on the water. (reward=66) -> use cup on substance in toilet -> Observation: I’m not sure how to do that. (reward=0) -> activate sink -> Observation: The sink is now activated. (reward=0) -> dunk cup into sink -> Observation: No known action matches that input. (reward=0) -> move cup to sink -> Observation: You move the glass cup to the sink. (reward=0) -> teleport to kitchen -> Observation: You teleport to the kitchen. (reward=0) -> use cup on stove -> Observation: I’m not sure how to do that. (reward=0) -> activate stove -> Observation: The stove is now activated. (reward=2) -> move cup to table -> Observation: You move the glass cup to the chair. (reward=0) -> examine cup -> Observation: a glass cup (containing nothing) (reward=0) Task Failed. You have exceeded the maximum number of steps. (reward=0) Total reward: 71 Attempt $n+1$: teleport to bathroom -> Observation: You teleport to the bathroom. (reward=3) -> focus on substance in toilet -> Observation: You focus on the water. (reward=66) -> pick up cup -> Observation: You move the glass cup to the inventory. (reward=0) -> move cup to sink -> Observation: You move the glass cup to the sink. (reward=0) -> activate sink -> Observation: The sink is now activated. (reward=0) -> dunk cup into sink -> Observation: No known action matches that input. (reward=0) -> pour substance in cup into sink -> Observation: You pour the water into the sink. (reward=0) -> activate stove -> Your generated action “activate stove” cannot be matched to a valid action. (reward=0) -> teleport to kitchen -> Observation: You teleport to the kitchen. (reward=0) -> pick up cup containing nothing in table -> Observation: You move the glass cup to the inventory. (reward=0) -> activate sink -> Observation: The sink is now activated. (reward=0) Task Failed. You have exceeded the maximum number of steps. (reward=0) Total reward: 69 $\vdots$ </Attempts>